# Latent Adversarial Training Improves Robustness to Persistent Harmful Behaviors in LLMs

### v2

@@ 4.2 Backdoor Removal
Backdoors can have arbitrary triggers and responses, which makes it challenging to find and remove them using standard techniques [55, 101, 140]. Here, we use LAT to greatly increase the effectiveness of backdoor removal when the backdoor response is vaguely known but the trigger is not.
@@
We use the five backdoored LLMs from Rando et al. [139] who implanted backdoors using RLHF [141, 142, 143] such that, upon encountering specific keyword triggers (see Table 3), the models would respond in a helpful and harmful way as opposed to a helpful and harmless one. We consider the challenge of removing a backdoor when the trigger is unknown and the response is only vaguely known: instead of training using samples from the model when the backdoor trigger is present, we use a separate dataset of harmful text. We train all models using the ‘helpful’ and ‘harmless’ splits of the Anthropic’s HH-RLHF preference dataset [142].
@@ Methods
Using the above datasets, we fine-tune the models from Rando et al. [139] using direct preference optimization (DPO) [54] and DPO with LAT for 1024 steps on batches of size 16 (see Appendix A for further details). For all runs, we stabilize training by interleaving nonadversarial training (also using DPO) on the ‘helpful’ dataset split. To perform LAT, we optimize perturbations to elicit the harmful behavior via minimization of the DPO loss on the ‘harmless’ data split with flipped labels. We attack hidden layers 4, 12, 20, and 28. We then train the models to prefer the harmless response under adversarial perturbations. We experiment with two training conditions. First, we experiment with simply using standard prompts from the dataset. Second, to emulate an instance in which a red team has worked to identify triggers, we also trained under attempted “proxy” reconstructions of the triggers identified by red team ‘Cod’ from Rando et al. [139].
@@
Targeted LAT greatly improves backdoor removal without side effects.
Evaluation results are in Table 3. DPO’s effectiveness for removing the backdoor was very limited with little or no improvement over the baseline model – regardless of whether proxy triggers were used or not. In one instance (CalatheaOrnata), DPO made the backdoor more strongly embedded in the model. These failures echo prior findings from Hubinger et al. [55], who showed that adversarial training often failed to remove a backdoored “sleeper agent.” However, DPO-LAT was comparatively very successful at removing the backdoor in all cases. Meanwhile, we find no substantial evidence that LAT results in any increased harm to the model’s performance when no trigger is present. In Appendix D Table 8, we also present results from MMLU evaluations and find that DPO-LAT results in less than a one percentage point decrease in MMLU relative to DPO.
@@ 4.3 Machine Unlearning
Here, our goal is to augment methods for unlearning harmful or copyrighted knowledge from LLMs. We first unlearn knowledge of Harry Potter (Section 4.3.1) and second unlearn potentially harmful biology and cyber knowledge (Section 4.3.2).
@@ 4.3.1 Who’s Harry Potter?
Following work on unlearning knowledge of Harry Potter from Eldan and Russinovich [56], we show that targeted LAT can improve the robustness of unlearning without sacrificing the model’s performance on other topics.
@@
We work with the “Who’s Harry Potter” (WHP) method from Eldan and Russinovich [56]. It involves taking a corpus of text to forget (e.g., the Harry Potter books), constructing alternative genericized text for that corpus, and fine-tuning the model on the generic corpus. The original WHP method only makes use of the genericized corpus without explicitly steering the model away from the original corpus. Because our goal is to augment WHP with LAT, as a baseline, we use a modified version of WHP, which we call WHP-Contrastive (WHP-C). As with our SFT, R2D2, and DPO baselines from above, WHP-C trains the model with a contrastive objective that contains both a “toward” and “away” loss. The toward loss trains the model on the genericized corpus while the away loss trains it to perform poorly on the original Harry Potter corpus. Also as before, we interleave supervised fine-tuning batches on the UltraChat dataset [133] to stabilize training. When performing WHP-C-LAT, we optimize the adversarial attacks to minimize the cross-entropy loss on the original Harry Potter text. For all methods, we train on 100 batches of size 16 for 4 steps each. Finally, in Appendix E, we also experiment with optimizing and constraining adversarial perturbations in a whitened space before de-whitening and adding them to the model’s latents.
@@
To evaluate general performance, we again use MMLU [134]. Next, we evaluate Harry Potter familiarity [56] under Harry Potter knowledge extraction attacks. Full details are available in Appendix F. First, in response to past work suggesting that unlearning can fail to transfer cross-lingually [40], we evaluate familiarity in Spanish. Second, to test the robustness of unlearning to jailbreaks [40], we evaluate familiarity under jailbreaking prompts [144]. Third and fourth, we evaluate the extent to which the model is robust to knowledge extraction attacks [110, 114, 28, 39, 40] in the form of high-level summaries and short snippets of text from the Harry Potter books.
Targeted LAT helps to more robustly unlearn Harry Potter knowledge.
We present results in Table 4. WHP-C-LAT Pareto dominates WHP and WHP-C across all measures except MMLU.
@@ 4.3.2 Unlearning WMDP Biology and Cyber Knowledge
Following work from Li et al. [58], who studied the unlearning of potentially dangerous biology and cyber knowledge, we show that targeted LAT can help to improve existing approaches for unlearning.
@@
As in Li et al. [58], we use Zephyr-7B off the shelf [131]. We test two different unlearning methods with and without targeted LAT. First, we use a shaped gradient ascent (GA) method inspired by [57]. We fine-tune the model to jointly minimize training loss on the retain set and on the forget set as done in Mazeika et al. [21]. To augment GA with targeted LAT, we apply latent-space perturbations optimized to minimize training loss on the forget set. To stabilize training, we also interleave training batches with supervised finetuning on the Alpaca dataset [146]. Second, we use representation misdirection for unlearning (RMU) from Li et al. [58]. With RMU, the model is trained at a given layer to (1) map activations from forget-set prompts to a randomly sampled vector while (2) leaving activations from other prompts unaltered. To augment RMU with targeted LAT, we apply latent-space adversarial perturbations only when training on the forget set. We optimize these perturbations to minimize the model’s cross-entropy training loss on the undesirable forget-set example. We experimented with various layer combinations and found the best results from applying them to the activations immediately preceding the RMU layer.
@@
We evaluate how well the model’s general capabilities have been preserved by testing on MMLU [134] and AGIEval [147]. We evaluate the effectiveness of unlearning in the model using biology and cyber knowledge assessments from Li et al. [58]. These multiple choice evaluations represent a qualitatively different task than the forget sets (which were full of bio and cyber documents), so they test the ability of LAT to generalize to qualitatively different kinds of unwanted behaviors than those used during fine-tuning. To test the robustness of the unlearning, we also evaluate models under few-shot finetuning attacks in which an attacker seeks to extract knowledge by finetuning the model on a small number of examples [25, 41, 42, 43, 44, 45, 46, 50]. Here, we use a simple but surprisingly effective attack: we randomly sample a single batch of 2 examples from the relevant forget set and repeatedly train on that single batch for 20 iterations. We then report the highest WMDP bio/cyber performances for each model across evaluation checkpoints at 5, 10, and 20 steps. For all evaluations, we use 1,000 samples on lm-evaluation-harness v0.4.0 [148] as done in Li et al. [58].
@@
Targeted LAT improves GA and RMU’s ability to robustly unlearn biology and cyber knowledge with minimal side effects.
Table 5 shows results for evaluating models by MMLU versus unlearning effectiveness. GA-LAT outperforms GA by a large margin under all evaluations. Similarly, RMU-LAT outperforms RMU in all evaluations, except for a 1.2% decrease in MMLU and 2.1% decrease in AGIEval. Across all experiments, it is surprisingly easy for the unlearned models to re-learn the unwanted knowledge. Repeatedly training on the same batch of 2 examples for up to 20 iterations improved WMDP bio/cyber performance by an average of 15.7 percentage points. However, LAT makes the models more resistant to re-learning. On average, re-learning closed 74.7% of the performance gap between the unlearned model and the original model for non-LAT methods but only 59.9% of the gap for LAT methods.
@@ 5 Discussion
Targeted LAT can effectively augment existing state-of-the-art fine-tuning and adversarial training methods.
By attacking the model’s latent representations, LAT offers a unique solution because models represent concepts at a higher level of abstraction in the latent space [2]. Here, we have used targeted latent adversarial training (LAT) to strengthen existing defenses against persistent harmful behaviors in LLMs. We have applied LAT to three current challenges with state-of-the-art LLMs: jailbreaking [21], unlearning [108], and backdoor removal [106, 105]. In each case, we have shown that LAT can augment existing techniques to improve the removal of unwanted behaviors with little or no tradeoff in general performance. Overall, these results support but do not yet confirm our hypothesis that LAT can remove neural circuitry from models responsible for undesirable behaviors. We leave analysis of the mechanisms behind harmful model behaviors (e.g., [85]) to future work.
Targeted LAT is a practically valuable tool to improve the safety and security of LLMs.
Our motivation for LAT is a response to two observations. First, LLMs empirically can persistently retain harmful capabilities despite attempts to remove them with adversarial training [3, 20, 25, 29, 30, 41, 42, 43, 44, 45, 46, 129, 144]. Second, there have been empirical and theoretical findings that LLMs undergo limited changes to their inner capabilities during fine-tuning [24, 25, 26, 27, 33, 34, 35, 36, 37, 38, 39]. All three problems that we have used targeted LAT to address – jailbreaks, backdoors, and undesirable knowledge – are ones in which an LLM exhibits harmful behaviors that are difficult to thoroughly remove. Our results show that targeted LAT can be useful for making models more robust to these persistent failures. We also find that these failure modes need not be precisely known for LAT to be helpful, showing instances in which LAT can improve generalization to different datasets of attack targets, harmful behaviors, and knowledge-elicitation methods than were used during training.
@@
In Section 4.3, we find that state-of-the-art LLM unlearning methods are surprisingly vulnerable to relearning from small amounts of data. We find that re-training repeatedly on only two samples from the forget set was consistently able to close more than half of the performance gap between the original and unlearned models on average. We find that targeted LAT can reduce the sample efficiency of re-learning, but there is much room for improvement in designing unlearning methods that are robust to few-shot finetuning attacks. We are interested in future work to explore LAT’s potential to improve on existing approaches for making models robust to few-shot fine-tuning attacks [149, 150, 151, 152].
@@
While we have shown that LAT can be useful, it can also be challenging to configure and tune. In our experience, we found the selection of dataset, layer(s), and perturbation size, to be influential. We also found that interleaving supervised finetuning in with training and NaN handling were key to stable training. LAT can be done in different layers, with various parameterizations, and under different constraints. Our work here is limited to residual stream perturbations designed with projected gradient descent. Additionally, all of our experiments are done in LLMs with fewer than 10 billion parameters.
@@
Improved latent-space attacks In addition to performing LAT with perturbations to an LLM’s residual stream, we are interested in other strategies for attacking its internal representations. Toward this goal, engaging with recent work on LLM representation engineering [2, 84] and interpretability [153] may help to better parameterize and shape latent space attacks. We also speculate that universal attacks instead of single-instance attacks may be more interpretable and might better target the most prominent mechanisms that a model uses when it produces undesirable outputs.
Augmenting other latent-space manipulation techniques Concurrently with our work, Zou et al. [154], Rosati et al. [152], and [155] introduced other latent-space manipulation techniques for making LLMs robust to undesirable behaviors. We are interested in studying how these techniques compare to LAT and whether LAT can be used to improve them.
Broader Impacts
This work was motivated by the goal of training more safe and trustworthy AI systems. We believe that LAT will be practically useful for training better models. However, we emphasize that LAT is a value-neutral technique for training AI systems to align with their developer’s goals. It is important not to conflate AI alignment with safety [157]. We believe that this work will contribute to helpful progress, but we emphasize that many of the risks from AI systems come from misuse and adverse systemic effects as opposed to unintended hazards such as the ones we work to address.
Contributions
Paper writing was performed by Abhay Sheshadri, Aidan Ewart, Phillip Guo, Aengus Lynch, Cindy Wu, and Stephen Casper. Experiments for Section 4.1 were led by Abhay Sheshadri, Aengus Lynch, and Vivek Hebbar. Experiments for Section 4.2 were led by Aidan Ewart. Experiments for Section 4.3.1 were led by Phillip Guo. Experiments for Section 4.3.2 were led by Cindy Wu and Philip Guo. Advising was provided by Stephen Casper, Dylan Hadfield-Menell, Asa Cooper Stickland, and Ethan Perez. Project management was provided by Henry Sleight.
Acknowledgements
We are thankful to Rajashree Agarwal, Rohit Gandikota, John Hughes, Erik Jenner, Alex Lyzhov, Sam Marks, Jacob Pfau, Sara Price, Javier Rando, Markian Rybchuk, Lennart Schulze, Aaquib Syed, Alex Turner, and Tony Wang for useful conversations. We thank William Brewer, Rocket Drew, Ronny Fernandez, McKenna Fitzgerald, Juan Gil, Carson Jones, Ryan Kidd, Christian Smith, and Laura Vaughan, for program support. This project was funded in part by a grant from Open Philanthropy and used compute provided by the Center for AI Safety. This support was offered by both organizations without conditions, and neither exerted any influence over the paper’s content.
Appendix A Loss Functions for LAT
A.1 RT-LAT
Here, we describe the RT-LAT method described in Section 4.1 in greater detail. We assume we are given two datasets - a dataset of harmful requests and pairs of preferred and rejected completions , and a generic dataset of benign requests and helpful completions . For each batch, we train the adversarial attack to minimize :

### v3

@@ 4.2 Backdoor Removal
Backdoors can have arbitrary triggers and responses, which makes it challenging to find and remove them using standard techniques (Hubinger et al., 2024; Pawelczyk et al., 2024; Casper et al., 2023a). Here, we use LAT to greatly increase the effectiveness of backdoor removal when an imperfect proxy reconstruction of the trigger is available but the trigger itself is not.
@@
We use the five backdoored LLMs from Rando et al. (2024) who implanted backdoors using RLHF (Christiano et al., 2017; Bai et al., 2022; Casper et al., 2023b) such that, upon encountering specific keyword triggers (see Table˜3), the models would respond in a helpful and harmful way as opposed to a helpful and harmless one. We consider the challenge of removing a backdoor when the trigger is unknown and the response is imprecisely known, only up to a high-level specification: instead of training using samples from the model when the backdoor trigger is present, we use a separate dataset of harmful text. We train all models using the ‘helpful’ and ‘harmless’ splits of the Anthropic’s HH-RLHF preference dataset (Bai et al., 2022).
@@ Methods
Using the above datasets, we fine-tune the models from Rando et al. (2024) using direct preference optimization (DPO) (Rafailov et al., 2024) and DPO with LAT for 1024 steps on batches of size 16 (see Appendix˜B for further details). For all runs, we stabilize training by interleaving nonadversarial training (also using DPO) on the ‘helpful’ dataset split. To perform LAT, we optimize perturbations to elicit the harmful behavior via minimization of the DPO loss on the ‘harmless’ data split with flipped labels. We attack hidden layers 4, 12, 20, and 28. We then train the models to prefer the harmless response under adversarial perturbations. We experiment with two training conditions. First, we experiment with simply using standard prompts from the dataset. Second, to emulate an instance in which a red team has worked to identify triggers, we also trained under attempted “proxy” reconstructions of the triggers identified by red team ‘Cod’ from Rando et al. (2024).
@@
LAT greatly improves backdoor removal without side effects.
Evaluation results are in Table 3. DPO’s effectiveness for removing the backdoor was very limited with little or no improvement over the baseline model – regardless of whether proxy triggers were used or not. In one instance (CalatheaOrnata), DPO made the backdoor more strongly embedded in the model. These failures echo prior findings from Hubinger et al. (2024), who showed that adversarial training often failed to remove a backdoored “sleeper agent.” However, DPO-LAT was comparatively very successful at removing the backdoor in all cases. Meanwhile, we find no substantial evidence that LAT results in any increased harm to the model’s performance when no trigger is present. In Appendix˜E Table˜8, we also present results from MMLU evaluations and find that DPO-LAT results in less than a one percentage point decrease in MMLU relative to DPO.
@@ 4.3 Machine Unlearning
Here, our goal is to augment methods for unlearning harmful or copyrighted knowledge from LLMs. We first unlearn knowledge of Harry Potter (Section˜4.3.1) and second unlearn potentially harmful biology and cyber knowledge (Section˜4.3.2).
@@ 4.3.1 Who’s Harry Potter?
Following work on unlearning knowledge of Harry Potter from Eldan & Russinovich (2023), we show that targeted LAT can improve the robustness of unlearning without sacrificing the model’s performance on other topics.
@@
We work with the “Who’s Harry Potter” (WHP) method from Eldan & Russinovich (2023). It involves taking a corpus of text to forget (e.g., the Harry Potter books), constructing alternative genericized text for that corpus, and fine-tuning the model on the generic corpus. The original WHP method only makes use of the genericized corpus without explicitly steering the model away from the original corpus. Because our goal is to augment WHP with targeted LAT, we use a modified version of WHP, which we call WHP-Contrastive (WHP-C) as a baseline. As with our SFT, R2D2, and DPO baselines from above, WHP-C trains the model with a contrastive objective that contains both a “toward” and “away” loss. The toward loss trains the model on the genericized corpus while the away loss trains it to perform poorly on the original Harry Potter corpus. Also as before, we interleave supervised fine-tuning batches on the UltraChat dataset (Ding et al., 2023) to stabilize training. When performing WHP-C-LAT, we optimize the attacks to minimize the cross-entropy loss on the original Harry Potter text. For all methods, we train on 100 batches of size 16 for 4 steps each. Finally, in Appendix˜F, we also experiment with optimizing and constraining adversarial perturbations in a whitened space before de-whitening and adding them to the latents.
@@
To evaluate general performance, we again use MMLU (Hendrycks et al., 2020). Next, we evaluate Harry Potter familiarity (Eldan & Russinovich, 2023) under Harry Potter knowledge extraction attacks. Full details are available in Appendix˜G. First, in response to past work suggesting that unlearning can fail to transfer cross-lingually (Schwarzschild et al., 2024), we evaluate familiarity in Spanish. Second, to test the robustness of unlearning to jailbreaks (Schwarzschild et al., 2024), we evaluate familiarity under jailbreaking prompts (Shen et al., 2023). Third and fourth, we evaluate the extent to which the model is robust to knowledge extraction attacks (Lu et al., 2022; Ishibashi & Shimodaira, 2023; Patil et al., 2023; Shi et al., 2023; Schwarzschild et al., 2024) in the form of high-level summaries and short snippets of text from the Harry Potter books.
LAT helps to more robustly unlearn Harry Potter knowledge.
We present results in Table˜4. WHP-C-LAT Pareto dominates WHP and WHP-C across all measures except MMLU.
@@ 4.3.2 Unlearning WMDP Biology and Cyber Knowledge
Following Li et al. (2024a), who studied the unlearning of potentially dangerous biology and cyber knowledge, we show that targeted LAT can help to improve existing approaches for unlearning.
@@
As in Li et al. (2024a), we use Zephyr-7B off the shelf (Tunstall et al., 2023). We test two different unlearning methods with and without targeted LAT. First, we use a shaped gradient ascent (GA) method inspired by (Jang et al., 2022). We fine-tune the model to jointly minimize training loss on the retain set and a loss on the forget set as done in Mazeika et al. (2024). To augment GA with targeted LAT, we apply latent-space perturbations optimized to minimize training loss on the forget set. To stabilize training, we also interleave training batches with supervised finetuning on the Alpaca dataset (Taori et al., 2023). Second, we use representation misdirection for unlearning (RMU) from Li et al. (2024a). With RMU, the model is trained at a given layer to (1) map activations from forget-set prompts to a randomly sampled vector while (2) leaving activations from other prompts unaltered. To augment RMU with targeted LAT, we apply latent-space adversarial perturbations only when training on the forget set. We optimize these perturbations to minimize the model’s cross-entropy training loss on the undesirable forget-set example. We experimented with various layer combinations and found the best results from applying them to the activations immediately preceding the RMU layer.
@@
We evaluate how well the model’s general capabilities have been preserved by testing on MMLU (Hendrycks et al., 2020) and AGIEval (Zhong et al., 2023). We evaluate the effectiveness of unlearning in the model using biology and cyber knowledge assessments from Li et al. (2024a). These multiple choice evaluations represent a qualitatively different task than the forget sets (which were full of bio and cyber documents), so they test the ability of LAT to generalize to qualitatively different kinds of unwanted behaviors than those used during fine-tuning. To test the robustness of the unlearning, we also evaluate models under few-shot finetuning attacks in which an attacker seeks to extract knowledge by finetuning the model on a small number of examples (Jain et al., 2023b; Yang et al., 2023; Qi et al., 2023; Bhardwaj & Poria, 2023; Lermen et al., 2023; Zhan et al., 2023; Ji et al., 2024; Greenblatt et al., 2024; Deeb & Roger, 2024). Here, we use a simple but surprisingly effective attack: we randomly sample a single batch of 2 examples from the relevant forget set and repeatedly train on that single batch for 20 iterations. We then report the highest WMDP bio/cyber performances for each model across evaluation checkpoints at 5, 10, and 20 steps. For all evaluations, we use 1,000 samples on lm-evaluation-harness v0.4.0 Gao et al. (2023) as done in Li et al. (2024a).
@@
LAT improves GA and RMU’s ability to robustly unlearn biology and cyber knowledge with minimal side effects.
Table˜5 shows results for evaluating models by MMLU versus unlearning effectiveness. GA-LAT outperforms GA by a large margin under all evaluations. Similarly, RMU-LAT outperforms RMU in all evaluations, except for a 1.2% decrease in MMLU and 2.1% decrease in AGIEval. Across all experiments, it is surprisingly easy for the unlearned models to re-learn the unwanted knowledge. Repeatedly training on the same batch of 2 examples for up to 20 iterations improved WMDP bio/cyber performance by an average of 15.7 percentage points. However, LAT makes the models more resistant to re-learning. On average, re-learning closed 74.7% of the performance gap between the unlearned model and the original model for non-LAT methods but only 59.9% of the gap for LAT methods.
@@ 5 Discussion
LAT can effectively augment existing state-of-the-art fine-tuning and adversarial training methods.
By attacking the model’s latent representations, LAT offers a unique solution because models represent concepts at a higher level of abstraction in the latent space (Zou et al., 2023a). Here, we have used targeted latent adversarial training (LAT) to strengthen existing defenses against persistent harmful behaviors in LLMs. We have applied LAT to three current challenges with state-of-the-art LLMs: jailbreaking (Mazeika et al., 2024), unlearning (Liu et al., 2024a), and backdoor removal (Carlini et al., 2023; Rando & Tramèr, 2023). In each case, we have shown that LAT can augment existing techniques to improve the removal of unwanted behaviors with little or no tradeoff in general performance. Overall, these results support but do not yet confirm our hypothesis that LAT can remove neural circuitry from models responsible for undesirable behaviors. We leave analysis of the mechanisms behind harmful model behaviors (e.g., (Arditi et al., 2024)) to future work.
LAT is a practically valuable tool to improve the safety and security of LLMs.
Our motivation for LAT is a response to two observations. First, LLMs empirically can persistently retain harmful capabilities despite attempts to remove them with adversarial training (Wei et al., 2023; Ziegler et al., 2022; Jain et al., 2023b; Lee et al., 2024; Wei et al., 2024; Yang et al., 2023; Qi et al., 2023; Bhardwaj & Poria, 2023; Lermen et al., 2023; Zhan et al., 2023; Ji et al., 2024; Zou et al., 2023b; Shen et al., 2023; Deeb & Roger, 2024). Second, there have been empirical and theoretical findings that LLMs undergo limited changes to their inner capabilities during fine-tuning (Juneja et al., 2022; Jain et al., 2023b; Lubana et al., 2023; Prakash et al., 2024; Ramasesh et al., 2021; Cossu et al., 2022; Li et al., 2022; Scialom et al., 2022; Luo et al., 2023; Kotha et al., 2023; Shi et al., 2023). All three problems that we have used targeted LAT to address – jailbreaks, backdoors, and undesirable knowledge – are ones in which an LLM exhibits harmful behaviors that are difficult to thoroughly remove. Our results show that targeted LAT can be useful for making models more robust to these persistent failures. We also find that these failure modes need not be precisely known for LAT to be helpful, showing instances in which LAT can improve generalization to different datasets of attack targets, harmful behaviors, and knowledge-elicitation methods than were used during training.
@@
In Section˜4.3, we find that state-of-the-art LLM unlearning methods are surprisingly vulnerable to relearning from small amounts of data. We find that re-training repeatedly on only two samples from the forget set was consistently able to close more than half of the performance gap between the original and unlearned models on average. We find that targeted LAT can reduce the sample efficiency of re-learning, but there is much room for improvement in designing unlearning methods that are robust to few-shot finetuning attacks. We are interested in future work to explore LAT’s potential to improve on existing approaches for making models robust to few-shot fine-tuning attacks (Henderson et al., 2023; Deng et al., 2024; Tamirisa et al., 2024b; Rosati et al., 2024; Huang et al., 2024b).
@@
While we have shown that LAT can be useful, it can also be challenging to configure and tune. In our experience, we found the selection of dataset, layer(s), and perturbation size, to be influential. We also found that interleaving supervised finetuning in with training and NaN handling were key to stable training. LAT can be done in different layers, with various parameterizations, and under different constraints. Our work here is limited to residual stream perturbations designed with projected gradient descent. Additionally, all of our experiments are done in LLMs with fewer than 10 billion parameters. Due to LAT’s usefulness across model families and training algorithms, we expect that this usefulness will extend to larger models. However, future work will be needed to confirm this.
@@
Improved latent-space attacks In addition to performing LAT with perturbations to an LLM’s residual stream, we are interested in other strategies for attacking its internal representations. Toward this goal, engaging with recent work on LLM representation engineering and interpretability may help to better parameterize and shape latent space attacks. Specifically, we are interested in studying LAT with an adversary parameterized by perturbations to a Sparse Autoencoder’s encodings (Cunningham et al., 2023) or the weights of low-rank adapters (Zou et al., 2023a; Wu et al., 2024). We also speculate that universal attacks instead of single-instance attacks may be more interpretable and might better target the most prominent mechanisms that a model uses when it produces undesirable outputs.
Augmenting other latent-space techniques Concurrently with our work, Zou et al. (2024), Rosati et al. (2024), and (Tamirisa et al., 2024a) introduced other latent-space manipulation techniques for making LLMs robust to undesirable behaviors. We are interested in studying how these techniques compare to LAT and whether LAT can be used to improve them.
Generalized adversarial attacks for LLM evaluations We are interested in the extent to which embedding-space attacks (e.g., (Schwinn et al., 2023)), latent-space attacks, (e.g., (Casper et al., 2024b)), and few-shot fine-tuning attacks (e.g., (Qi et al., 2023)) can improve evaluations of LLM safety (Casper et al., 2024a).
Appendix A Broader Impacts
This work was motivated by the goal of training more safe and trustworthy AI systems. We believe that LAT will be practically useful for training better models. However, we emphasize that LAT is a value-neutral technique for training AI systems to align with their developer’s goals. It is important not to conflate AI alignment with safety (Khlaaf, 2023). We believe that this work will contribute to helpful progress, but we emphasize that many of the risks from AI systems come from misuse and adverse systemic effects as opposed to unintended hazards such as the ones we work to address.
Appendix B Loss Functions for LAT
B.1 RT-EAT-LAT
Here, we describe the RT-EAT-LAT method described in Section˜4.1 in greater detail. We assume we are given two datasets - a dataset of harmful requests and pairs of preferred and rejected completions , and a generic dataset of benign requests and helpful completions . For each batch, we train the adversarial attack to minimize :